The following data are available for two divisions of Solomons Company.

North Division South Division
Division operating profit $ 6,720,000 $ 43,470,000
Division investment 32,000,000 322,000,000

The cost of capital for the company is 7 percent. Ignore taxes.  

Required:

a-1. Calculate the ROI for both North and South divisions.

a-2. If Solomons measures performance using ROI, which division had the better performance?

b-1. Calculate the EVA for both North and South divisions. (The divisions have no current liabilities.)

b-2. If Solomons measures performance using economic value added, which division had the better performance?

c. Would your evaluation change if the company’s cost of capital was 16 percent?

1. When evaluated by ROI?

2. When evaluated by EVA?
Answer #1

Calculation of ROI for both the divisions, {Operating profit/investment}

  1. North division,
    • 6,720,000÷32,000,000 = 21%
  2. South division,
    • 43,470,000÷322,000,000 = 13.5%

If Solomon's measures performance using ROI then north division has a better ROI than south division.

Calculation of EVA, {NOPAT - (WACC*capital invested)

  1. North division,
    • 6,720,000 - (32,000,000*7%) = 4,480,000.
  2. South division,
    • 43,470,000 - (322,000,000*7%) = 20,930,000.

Under EVA, the south division is performing better.

When Cost of capital is 16%,

ROI, it won't be affected if the cost of capital changes, but the thing is that south division doesn't even meet the cost of capital of 16%.

EVA,

  1. North division,
    • 6,720,000 - (32,000,000*16%) = 1,600,000.
  2. South division,
    • 43,470,000 - (322,000,000*16%) = 8,050,000.

Here also, south division is performing better,

Evaluation remains the same, no matter the cost of capital is 7% or 16%.